OpenStudy (anonymous):

Use L'Hôpital's rule to solve these questions: I'm going to put them in a second

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ x cosx - sinx }{ xsin ^{2}x }\]

OpenStudy (anonymous):

this is the first question , I already know the answer , I want someone to explain , because in my notes " from my prof " he said that the answer is -1/3 and the final step is going to be: \[\frac{ -1 }{ 2x sinx+3cosx }\] and then we find the limit to get -1/3 but when I tried to solve it , I got : \[\frac{ 1 }{ 2+sinx cosx }\] and I got 1/3 as the answer.

OpenStudy (anonymous):

the second questions : also using the same rule to find : \[\lim_{x \rightarrow 0+} xLnx\]

OpenStudy (anonymous):

about the first question , the last step that I did is going to be : \[\frac{ 1 }{ 2+\cos ^{^{2}}x } then find the limit\]